OpenStudy (anonymous):

f(x) = (2x+1)^(1/2) g is the inverse of f(x) g'(3) = ?

OpenStudy (kirbykirby):

You can say \(f(x)\) is \(y\), and then swap \(x\) and \(y\) to find the inverse: \[ (2y+1)^{1/2}=x\\ \left((2y+1)^{1/2}\right)^2= x^2\\ (2y+1)=x^2\\ 2y=x^2-1\\ y=\frac{x^2-1}{2}\] Now, this is your \(g(x)\) function, i.e. \(g(x)=\frac{x^2-1}{2}\). Now you need to find \(g'(x)\) \[g'(x)=\frac{1}{2}(2x)=x\] Then, \(g'(3)=3\)